The paths of two ships are tracked on the same coordinate system. One ship is following a path described by the equation 2x + 3y = 6 , and the other is following a path described by the equation y = 2/3x - 6 . What are the coordinates of the collision point?
Enter your answer as an ordered pair.

Answer by josmiceli(19441)

Answer is the coordinates that solve both equations.
2x + 3y = 6
2/3x -y = 6 (rearraging)
multiply both sides of the 2st equation by 3.
2x + 3y = 6
2x - 3y = 18
add the equations
4x = 24
divide by 4
x = 6
plug this into 1st equation
2*6 + 3y = 6
12 + 3y = 6
3y = -6
y = -2
(6, -2) is the collision point
check it
2/3*6 -(-2) = 6
4 +2 = 6
that checks
2*6 + 3(-2) = 6
12 - 6 = 6
that checks also